# Communication-Efficient and Privacy-Adaptable Mechanism for Federated Learning

### v1

@@ II-C Differential Privacy (DP)
The group privacy profile $\delta_{\mathcal{F},j}(\tilde{\epsilon})$ ($j\geq 1$) is defined as:
$\delta_{\mathcal{F},j}(\tilde{\epsilon}):=\sup_{d(X,X^{\prime})\leq k}D_{e^{\epsilon}}(\mathcal{F}(X)||\mathcal{F}(X^{\prime})),$ (11)
@@
In [38], Balle et al. demonstrated a result on privacy amplification with subsampling with replacement as follows.
Theorem 7
[38, Theorem 10] Let $s$ be the subsampling with replacement mechanism and let $\mathcal{F}:\mathcal{X}\rightarrow\mathcal{Y}$ be a base privacy mechanism. Then, for $\tilde{\epsilon}\geq 0$ and set $\epsilon=\log\left(1+p(e^{\tilde{\epsilon}}-1)\right)$ where $p:=1-\left(1-\frac{1}{n}\right)^{\tau}$, we have
$\delta_{\mathcal{F}(s(X))}(\epsilon)\leq\sum_{j=1}^{\tau}\binom{\tau}{j}\left(\frac{1}{n}\right)^{j}\left(1-\frac{1}{n}\right)^{\tau-j}\delta_{\mathcal{F}(X),j}(\tilde{\epsilon}),$ (12)
@@
The above result guarantees that utilizing a differentially private mechanism on a random subsample of a dataset in a single communication round of FL framework offers stronger privacy assurances compared to applying it on the whole dataset.
@@
Theorem 8
[56, Theorem 8] Let $q:\mathcal{X}\rightarrow\mathbb{R}^{n}$ be a function with global $\ell_{2}$-sensitivity $\Delta_{2}:=\sup_{(X,X^{\prime})\in\mathcal{R}}\|q(X)-q(X^{\prime})\|_{2}$. Then, for any $\epsilon\geq 0$ and $\delta\in[0,1]$, the Gaussian mechanism $\mathcal{F}(X):=q(X)+\mathbf{Z}$ with $\mathbf{Z}\sim\mathcal{N}(\mathbf{0},\tilde{\sigma}^{2}\mathbf{I}_{n})$, where $\mathbf{I}_{n}$ is the $n\times n$ identity matrix, is $(\epsilon,\delta)$-DP if and only if
@@
Lemma 9
Let $q:\mathcal{X}\rightarrow\mathbb{R}^{n}$ be a function with global $\ell_{2}$-sensitivity $\Delta_{2}:=\sup_{(X,X^{\prime})\in\mathcal{R}}\|q(X)-q(X^{\prime})\|_{2}$, and let $s$ be the subsampling with replacement mechanism. Then, for any $\epsilon\geq 0$ and $\delta\in[0,1]$, the subsampled Gaussian mechanism $M(X)=q\circ s(X)+N(\mathbf{0},\tilde{\sigma}^{2}\mathbf{I}_{n})$ is $(\epsilon,\delta)$-DP if and only if, for any $\tilde{\epsilon}>0$, $\epsilon=\log\left(1+p(e^{\tilde{\epsilon}}-1)\right)$ where $p:=1-\left(1-\frac{1}{n}\right)^{\tau}$ and
@@
$\delta_{\mathcal{F}(X),j}(\tilde{\epsilon})\leq\frac{(e^{\tilde{\epsilon}}-1)\delta_{\mathcal{F}(X)}(\tilde{\epsilon}/j)}{e^{\tilde{\epsilon}/j}-1}.$
@@
$\sum_{j=1}^{\tau}\binom{\tau}{j}\left(\frac{1}{n}\right)^{j}\left(1-\frac{1}{n}\right)^{\tau-j}\frac{(e^{\tilde{\epsilon}}-1)\delta_{\mathcal{F}(X)}(\tilde{\epsilon}/j)}{e^{\tilde{\epsilon}/j}-1}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{j=1}^{\tau}\binom{\tau}{j}\left(\frac{1}{n}\right)^{j}\left(1-\frac{1}{n}\right)^{\tau-j}\frac{(e^{\tilde{\epsilon}}-1)}{e^{\tilde{\epsilon}/j}-1}\left(\Phi\left(\frac{\Delta_{2}}{2\tilde{\sigma}}-\frac{\tilde{\epsilon}\tilde{\sigma}}{j\Delta_{2}}\right)-e^{\tilde{\epsilon}/j}\Phi\left(-\frac{\Delta_{2}}{2\tilde{\sigma}}-\frac{\tilde{\epsilon}\tilde{\sigma}}{j\Delta_{2}}\right)\right),$
@@
Theorem 10
[38, Theorem 3] Let $q:\mathcal{X}\rightarrow\mathbb{R}$ be a function with the global $\ell_{1}$-sensitivity $\Delta_{1}:=\sup_{(X,X^{\prime})\in\mathcal{R}}|q(X)-q(X^{\prime})|$ and $\mathcal{F}(X):=q(X)+Z$ with $Z\sim\mathrm{Lap}(0,b)$ be the Laplace mechanism. Then, for any $\epsilon>0$, the privacy profile of $\mathcal{F}$ is given by $\delta_{\mathcal{F}}(\epsilon)=\left[1-e^{\frac{\epsilon-\Delta_{1}/b}{2}}\right]_{+}$, where $[\cdot]_{+}:=\max\{\cdot,0\}$.
@@
II-D Problem Setting
Our goal is to construct a mechanism that simultaneously addresses privacy requirements (protecting against honest but inquisitive clients within the FL network or inquisitive data analysts upon public release of the final trained model, see Section II-D1) and compression demands (for lossless uplink channels with limited bandwidth) in a single local model update at each client within the FL framework. Since the distribution of the model parameters and/or the induced gradient is often unknown to the clients, our focus lies in creating a universal mechanism applicable to any random source. Furthermore, we are exploring privacy mechanisms that introduce noise customized to various noise distributions. While Laplace mechanism provides pure DP protection, Gaussian mechanism only provides approximate DP protection with a small failure probability [10, 11]. However, it is well-known that Gaussian mechanism support tractability of the privacy budget in mean estimation [58, 54], an important subroutine in FL. Such schemes can be formulated as mappings from the local update $\mathbf{X}_{t}^{k}\in\mathbb{R}^{m}$ at client $k$ at the time instance $t$ to the estimated update $\hat{\mathbf{X}}_{t}^{k}\in\mathbb{R}^{m}$ at the server, aimed to achieve the following desired properties:
@@ II-D1 Threat model
III CEPAM for FL
In this section, we introduce CEPAM, which utilizes the inherent randomness in probabilistic quantized FL to provide privacy enhancement. (see Algorithm 1). Our approach modifies the schemes in [9] and [22] by replacing universal quantization [46, 47] with RSUQ as outlined in Section II-B. Through the utilization of RSUQ and the layered construction [35, 34], leveraging on the unique quantization error property of RSUQ (refer to Proposition 5), we construct a joint privacy and compression mechanism to address both FL challenges of communication overload and privacy requirement adaptable to any system adjustment simultaneously. This is achieved by applying LRSUQ on the model updates $\mathbf{X}_{t+\tau}^{k}$ at the end of each FL round for every client separately. These quantized updates are then transmitted as a set of messages $\{\mathbf{M}_{j}^{k}\}_{j\in\mathcal{N}}$ to the server. The server adds the dithers induced by the shared randomness to the received messages, collects the decoded messages and obtains the estimations $\hat{\mathbf{X}}_{t}^{k}$. Subsequently, the server computes the average of estimated model updates $\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t}^{k}$, which ensures the DP guarantee against the honest but inquisitive clients. The performance of CEPAM will be analyzed in Section IV.
@@ III CEPAM for FL
1:Inputs: Number of total iterations $T$, number of local iterations $\tau$, number of clients $K$, local datasets $\{\mathcal{D}^{(k)}\}_{k\in\mathcal{K}}$, loss function $\ell(\cdot,\cdot)$
@@
3:Initialization: Client $k$ and the server agree on privacy budget $\epsilon>0$ and privacy relaxation $\delta$ for $(\epsilon,\delta)$-DP (or privacy budget $\epsilon>0$ for $\epsilon$-DP), shared seed $s_{k}$, lattice dimension $n$, generator matrix $\mathbf{G}$, scaling coefficient $\gamma>0$, privacy-preserving noise $\mathbf{Z}\sim f$ with noise variance $\mathrm{Var}(f)>0$, latent variable $U\sim g(u):=\mu(L_{u}^{+}(f))$, initial model parameter vector $\mathbf{W}_{0}$
@@
9:         Compute $\mathbf{W}_{t+t^{\prime}}^{k}\leftarrow\mathbf{W}_{t}^{k}-\eta_{t+t^{\prime}}\nabla F_{k}^{j_{t}^{k}}(\mathbf{W}_{t}^{k})$
@@
12:     Run subroutine $\mathrm{ENCODE}(\mathbf{X}_{t+\tau}^{k},\gamma,s_{k},(f,g),N)$triangleright See Algorithm 2
13:     Send $\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}}$ to server, using $N\cdot\left(H(\mathrm{Geom}(p(U)\;|\;U)+H(\lceil\log|\mathcal{M}(U)|\rceil\;|\;U)\right)$ bits
14:end for
15:Protocol at the server:
16:for $t+\tau\in\mathcal{T}_{T}$ do
17:     for $k\in\mathcal{K}$ do
18:         Run subroutine $\mathrm{DECODE}(\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}},\gamma,s_{k},g,N)$triangleright See Algorithm 3
19:     end for
20:     Compute $\hat{\mathbf{W}}_{t+\tau}\leftarrow\mathbf{W}_{t}+\sum_{k\in\mathcal{K}}p_{k}\hat{\mathbf{X}}_{t+\tau}^{k}$
21:     Set $\mathbf{W}_{t+\tau}\leftarrow\hat{\mathbf{W}}_{t+\tau}$ and broadcast $\mathbf{W}_{t+\tau}$ to all clients, or output $\mathbf{W}_{T}$ if $t+\tau=T$
22:end for
@@
1) Initialization: At the outset, client $k$ and the server agree on the privacy budget $\epsilon$ and privacy relaxation $\delta$ for $(\epsilon,\delta)$ DP (or $\epsilon$ for $\epsilon$-DP) in accordance with the system requirements, along with the parameters for LRSUQ as specified in Section II-B. The latter involves sharing a small random seed $s_{k}$ between client $k$ and the server to serve as a source of common randomness while fixing the lattice dimension $n$, a lattice generator matrix $\mathbf{G}$, and a scaling coefficient $\gamma>0$. Both client $k$ and the server use the same random seed to initialize their respective PRNGs $\mathfrak{P}$ and $\tilde{\mathfrak{P}}$, ensuring that the outputs of the two PRNGs remain identical.
@@
Algorithm 2 $\mathrm{ENCODE}(\mathbf{X}_{t+\tau}^{k},\gamma,s_{k},(f,g),N)$
1:Inputs: model update vector ${\mathbf{X}}_{t+\tau}^{k}$, scaling coefficient $\gamma$, random seed $s_{k}$, pair of pdfs $(f,g)$, number of sub-vectors $N$
@@
3:Compute $\tilde{\mathbf{X}}_{t+\tau}^{k}\leftarrow\mathbf{X}_{t+\tau}^{k}/(\gamma^{-1}\|\mathbf{X}_{t+\tau}^{k}\|)$triangleright Perform normalizing
4:Partition $\tilde{\mathbf{X}}_{t+\tau}^{k}$ to $\{\tilde{\mathbf{X}}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$
5:for $j=1$ to $N$ do
6:     Sample $U_{t+\tau,j}^{k}\sim g$ by $\mathfrak{B}^{k}$triangleright Initiated by seed $s_{k}$
7:     for $i=1,2,\ldots$ do $\triangleright$ Perform RSUQ
8:          Sample $\mathbf{V}_{t+\tau,j,i}^{k}\sim\mathrm{Unif}(\mathcal{P})$ by $\mathfrak{B}^{k}$
9:         Find unique $\mathbf{M}_{t+\tau,j}^{k}\leftarrow Q_{\mathcal{P}}(\tilde{\mathbf{X}}_{t+\tau,j}^{k}/\beta(U_{t+\tau,j}^{k})-\mathbf{V}_{t+\tau,j,i}^{k})\in\mathbf{G}\mathbb{Z}^{n}$triangleright Perform encoding
10:          Check if $\beta(U_{t+\tau,j}^{k})\cdot\mathbf{M}_{t+\tau,j}^{k}-\tilde{\mathbf{X}}_{t+\tau,j}^{k}\in L_{U_{t+\tau,j}^{k}}^{+}(f)$
11:         if Yes then $\triangleright$ Perform rejection sampling
12:              $H_{t+\tau,j}^{k}\leftarrow i$; Return $(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})$
13:         else
14:              Reject $i$ and repeat Step 8-10 with $i+1$
15:         end if
@@
17:end for
18:return $\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}}$
Quantization: The $k$-th client normalizes the update $\mathbf{X}_{t+\tau}^{k}$ by a real number $\gamma^{-1}\|\mathbf{X}_{t+\tau}^{k}\|>0$ to $\tilde{\mathbf{X}}_{t+\tau}^{k}$, and then divides $\tilde{\mathbf{X}}_{t+\tau}^{k}$ into $N:=\left\lceil\frac{m}{n}\right\rceil$ distinct $n$-dimensional sub-vectors $\tilde{\mathbf{X}}_{t+\tau,j}^{k}\in\mathbb{R}^{n}$, $j\in\{1,\ldots,N\}=:\mathcal{N}$. The overhead associated with compressing the scalar $\gamma^{-1}\|\mathbf{X}_{t+\tau}^{k}\|$ is assumed to be negligible compared to compressing the vector $\tilde{\mathbf{X}}_{t+\tau}^{k}$.

### v2

@@ III-C Differential Privacy (DP)
In this work, we use the principle of privacy amplification by subsampling [39], whereby the privacy guarantees of a differentially private mechanism are amplified by applying it to a random subsample of the dataset. The problem of privacy amplification can be stated as follows. Let $\mathcal{F}:\mathcal{X}\rightarrow\mathcal{Y}$ be a privacy mechanism with privacy profile $\delta_{\mathcal{F}}$ with respect to the adjacent relation defined on $\mathcal{X}$, and let $s:\mathcal{W}\rightarrow\mathcal{X}$ be a subsampling mechanism. Consider the subsampled mechanism $\mathcal{F}^{s}:\mathcal{W}\rightarrow\mathcal{Y}$ given by $\mathcal{F}^{s}(X):=\mathcal{F}(s(X))$. The goal is to relate the privacy profile of $\mathcal{F}$ and $\mathcal{F}^{s}$.
@@
[…]
$\delta_{\mathcal{F},j}(\epsilon):=\sup_{d(X,X^{\prime})\leq k}D_{e^{\epsilon}}(\mathcal{F}(X)||\mathcal{F}(X^{\prime})),$ (11)
@@
We briefly review subsampling with replacement [39]. Let $\mathcal{U}$ be a set. We write $2^{\mathcal{U}}$ and $\mathbb{N}^{\mathcal{U}}$ for the collections of all sets and multisets, respectively, over $\mathcal{U}$. For an integer $n\geq 0$, we also write $2_{n}^{\mathcal{U}}$ and $\mathbb{N}_{n}^{\mathcal{U}}$ for the collections of all sets and multisets containing exactly $n$ elements, where the elements are counted with multiplicity for multisets. Given a multiset $X\in\mathbb{N}^{\mathcal{U}}$ we write $X_{a}$ for the number of occurrences of $a\in\mathcal{U}$ in $X$. The support of a multiset $X$ is the defined as the set $\mathrm{supp}(X):=\{a\in\mathcal{U}:X_{a}>0\}$. Given multisets $X,X^{\prime}\in\mathbb{N}^{\mathcal{U}}$ we write $X^{\prime}\subseteq X$ to denote that $X^{\prime}_{a}\leq X_{a}$ for all $a\in\mathcal{U}$. For order-independent datasets represented as multisets, the substitute-one relation, denoted as $X\simeq_{s}X^{\prime}$, holds whenever $\|X-X^{\prime}\|_{1}=2$ and $|X|=|X^{\prime}|$, i.e., $X^{\prime}$ is obtained by replacing an element in $X$ with a different element from $\mathcal{U}$. The subsampling with replacement mechanism $s_{\tau}:2_{n}^{\mathcal{U}}\rightarrow\mathfrak{M}(\mathbb{N}_{\tau}^{\mathcal{U}})$ takes a set of $x$ of size $n$ and outputs a sample from the multinomial distribution $\omega=s(X)$ over all multisets $Y$ of size $\tau\leq n$ and $\mathrm{supp}(Y)\subseteq X$. We further assume that the base mechanism $\mathcal{F}:\mathbb{N}_{\tau}^{\mathcal{U}}\rightarrow\mathcal{Y}$ is defined on the multisets and has privacy profile $\delta_{\mathcal{F}}$ with respect to $\simeq_{s}$. In [39], Balle et al. demonstrated a bound on the privacy profile of the subsampled mechanism with replacement $\mathcal{F}^{s_{m}}:2_{n}^{\mathcal{U}}\rightarrow\mathcal{Y}$ with respect to $\simeq_{s}$ as follows.
Theorem 7.
[39, Theorem 10] Let $s_{\tau}:2_{n}^{\mathcal{U}}\rightarrow\mathfrak{M}(\mathbb{N}_{\tau}^{\mathcal{U}})$ be the subsampling with replacement mechanism and let $\mathcal{F}:{\color[rgb]{0,0,0}\mathbb{N}_{\tau}^{\mathcal{U}}}\rightarrow\mathcal{Y}$ be a base privacy mechanism. Then, for $\tilde{\epsilon}\geq 0$ and $\epsilon=\log\left(1+p(e^{\tilde{\epsilon}}-1)\right)$ where $p:=1-\left(1-\frac{1}{n}\right)^{\tau}$, we have
[…]
@@
The above result guarantees that applying a differentially private mechanism to a random subsample of a dataset offers stronger privacy assurances than applying it to the entire dataset. This is because information about an individual cannot be leaked if that individual’s data is not included in the subsample.
@@
Theorem 8.
[71, Theorem 8] Let $q:\mathcal{X}\rightarrow\mathbb{R}^{n}$ be a function with global $\ell_{2}$-sensitivity $\Delta_{2}:=\sup_{(X,X^{\prime})\in\mathcal{R}}\|q(X)-q(X^{\prime})\|_{2}$. Then, for any $\epsilon\geq 0$ and $\delta\in[0,1]$, the Gaussian mechanism $\mathcal{F}(X):=q(X)+\mathbf{Z}$ with $\mathbf{Z}\sim\mathcal{N}(\mathbf{0},\tilde{\sigma}^{2}\mathbf{I}_{n})$, where $\mathbf{I}_{n}$ is the $n\times n$ identity matrix, is $(\epsilon,\delta)$-DP if and only if
@@
Lemma 9.
Let $q:\mathcal{X}\rightarrow\mathbb{R}^{n}$ be a function with global $\ell_{2}$-sensitivity $\Delta_{2}:=\sup_{(X,X^{\prime})\in\mathcal{R}}\|q(X)-q(X^{\prime})\|_{2}$, and let $s_{\tau}:2_{n}^{\mathcal{U}}\rightarrow\mathfrak{M}(\mathbb{N}_{\tau}^{\mathcal{U}})$ be the subsampling with replacement mechanism. Then, for any $\epsilon\geq 0$ and $\delta\in[0,1]$, the subsampled Gaussian mechanism ${\color[rgb]{0,0,0}\mathcal{F}^{s_{\tau}}(X)}=q\circ s(X)+N(\mathbf{0},\tilde{\sigma}^{2}\mathbf{I}_{n})$ is $(\epsilon,\delta)$-DP if and only if, for any $\tilde{\epsilon}>0$, $\epsilon=\log\left(1+p(e^{\tilde{\epsilon}}-1)\right)$ where $p:=1-\left(1-\frac{1}{n}\right)^{\tau}$ and
@@
$\delta_{\mathcal{F}(X),j}(\tilde{\epsilon})\leq\frac{(e^{\tilde{\epsilon}}-1)}{e^{\tilde{\epsilon}/j}-1}\cdot\delta_{\mathcal{F}(X)}(\tilde{\epsilon}/j).$
@@
$\sum_{j=1}^{\tau}\binom{\tau}{j}\left(\frac{1}{n}\right)^{j}\left(1-\frac{1}{n}\right)^{\tau-j}\frac{(e^{\tilde{\epsilon}}-1)}{e^{\tilde{\epsilon}/j}-1}\cdot\delta_{\mathcal{F}(X)}(\tilde{\epsilon}/j)$
$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{j=1}^{\tau}\binom{\tau}{j}\left(\frac{1}{n}\right)^{j}\left(1-\frac{1}{n}\right)^{\tau-j}\frac{(e^{\tilde{\epsilon}}-1)}{e^{\tilde{\epsilon}/j}-1}\left(\Phi\left(\frac{\Delta_{2}}{2\tilde{\sigma}}-\frac{\tilde{\epsilon}\tilde{\sigma}}{j\Delta_{2}}\right)-e^{\tilde{\epsilon}/j}\Phi\left(-\frac{\Delta_{2}}{2\tilde{\sigma}}-\frac{\tilde{\epsilon}\tilde{\sigma}}{j\Delta_{2}}\right)\right),$
@@
Lemma 9 implies that applying a subsampled Gaussian mechanism to a random subsample (obtained by sampling with replacement) of a dataset in a single communication round of the FL framework provides stronger privacy guarantees than applying the Gaussian mechanism to the entire dataset. Since each $k$-th client performs $\tau$ iterations of local SGD (i.e., Equation (3)) on indices sampled uniformly from the dataset $\mathcal{D}^{(k)}$ (i.e., $p_{1}=p_{2}=\ldots=p_{|\mathcal{D}^{(k)}|}=1/|\mathcal{D}^{(k)}|=1/n$, with $n=|\mathcal{D}^{(k)}|$ in (12) or (14), where $p_{j}$ denotes the probability that the data point $\xi_{k,j}$ is chosen at time instance $t$), this procedure is equivalent to subsampling with replacement, which outputs a multiset sample of size $\tau$ following a multinomial distribution $\mathrm{Mult}(\tau;p_{1},\ldots,p_{|\mathcal{D}^{(k)}|})$. [^6] See [21, Lemma 3] for an analogous result for the subsampled Gaussian mechanism under Poisson subsampling. The random multiset sample is then fed into the function $q(\cdot)=(\mathbf{W}_{t+\tau}^{k}-\mathbf{W}_{t})(\cdot)$. Refer to Theorem 12.
@@
Theorem 10.
[39, Theorem 3] Let $q:\mathcal{X}\rightarrow\mathbb{R}$ be a function with the global $\ell_{1}$-sensitivity $\Delta_{1}:=\sup_{(X,X^{\prime})\in\mathcal{R}}|q(X)-q(X^{\prime})|$ and $\mathcal{F}(X):=q(X)+Z$ with $Z\sim\mathrm{Lap}(0,b)$ be the Laplace mechanism. Then, for any $\epsilon>0$, the privacy profile of $\mathcal{F}$ is given by $\delta_{\mathcal{F}}(\epsilon)=\left[1-e^{\frac{\epsilon-\Delta_{1}/b}{2}}\right]_{+}$, where $\cdot]_{+}:=\max\{\cdot,0\}$.
@@
III-D Problem Setting
Our goal is to construct a mechanism that simultaneously addresses privacy requirements (protecting against honest but inquisitive clients within the FL network or inquisitive data analysts upon public release of the final trained model, see Section III-D1) and compression demands (for lossless uplink channels with limited bandwidth) in a single local model update at each client within the FL framework. Since the distribution of the model parameters and/or the induced gradient is often unknown to the clients, our focus lies in creating a universal mechanism applicable to any random source. Furthermore, we are exploring privacy mechanisms that introduce noise customized to various noise distributions. While Laplace mechanism provides pure DP protection, Gaussian mechanism only provides approximate DP protection with a small failure probability [10, 11]. However, it is well-known that Gaussian mechanism support tractability of the privacy budget in mean estimation [73, 69], an important subroutine in FL. Such schemes can be formulated as mappings from the local update $\mathbf{X}_{t}^{k}\in\mathbb{R}^{m}$ at client $k$ at the time instance $t$ to the estimated update $\hat{\mathbf{X}}_{t}^{k}\in\mathbb{R}^{m}$ at the server, aimed to achieve the following desired properties:
@@ III-D1 Threat model
IV CEPAM for FL
In this section, we introduce CEPAM, which utilizes the inherent randomness in probabilistic quantized FL to provide privacy enhancement. (see Algorithm 1). Our approach modifies the schemes in [9] and [21] by replacing universal quantization [57, 58] with RSUQ as outlined in Section III-B. Through the utilization of RSUQ and the layered construction [37, 36], leveraging on the unique quantization error property of RSUQ (refer to Proposition 5), we construct a joint privacy and compression mechanism to address both FL challenges of communication overload and privacy requirement adaptable to any system adjustment simultaneously. This is achieved by applying norm clipping and then LRSUQ to the model updates $\mathbf{X}_{t+\tau}^{k}$ at the end of each FL round, separately for each client. These quantized updates are then transmitted as a set of messages $\{\mathbf{M}_{j}^{k}\}_{j\in\mathcal{N}}$ to the server. The server adds the dithers induced by the shared randomness to the received messages, collects the decoded messages and obtains the estimations $\hat{\mathbf{X}}_{t}^{k}$. Subsequently, the server computes the average of estimated model updates $\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t}^{k}$, which ensures the DP guarantee against the honest but inquisitive clients. The performance of CEPAM will be analyzed in Section V.
@@ IV CEPAM for FL
1:Inputs: Number of total iterations $T$, number of local iterations $\tau$, number of clients $K$, local datasets $\{\mathcal{D}^{(k)}\}_{k\in\mathcal{K}}$, loss function $\ell(\cdot,\cdot)$, threshold $\gamma>0$
@@
3:Initialization: Client $k$ and the server agree on privacy budget $\epsilon>0$ and privacy relaxation $\delta$ for $(\epsilon,\delta)$-DP (or privacy budget $\epsilon>0$ for $\epsilon$-DP), shared seed $s_{k}$, lattice dimension $n$, generator matrix $\mathbf{G}$, privacy-preserving noise $\mathbf{Z}\sim f$ with noise variance $\mathrm{Var}(f)>0$, latent variable $U\sim g(u):=\mu(L_{u}^{+}(f))$, initial model parameter vector $\mathbf{W}_{0}$
@@
9:   Compute $\mathbf{W}_{t+t^{\prime}}^{k}\leftarrow\mathbf{W}_{t{\color[rgb]{0,0,0}+t^{\prime}-1}}^{k}-\eta_{t+t^{\prime}}\nabla F_{k}^{j_{t{\color[rgb]{0,0,0}+t^{\prime}-1}}^{k}}(\mathbf{W}_{t{\color[rgb]{0,0,0}+t^{\prime}-1}}^{k})$
@@
12:  Compute $\tilde{\mathbf{X}}_{t+\tau}^{k}\leftarrow\mathbf{X}_{t+\tau}^{k}/\max\{1,\|\mathbf{X}_{t+\tau}^{k}\|_{2}/\gamma\}$ $\triangleright$ Perform norm clipping
13:  Run subroutine $\mathrm{ENCODE}({\color[rgb]{0,0,0}\tilde{\mathbf{X}}_{t+\tau}^{k}},\gamma,s_{k},(f,g),N)$ $\triangleright$ See Algorithm 2
14:  Send $\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}}$ to server, using $N\cdot\left(H(\mathrm{Geom}(p(U)\;|\;U)+H(\lceil\log|\mathcal{M}(U)|\rceil\;|\;U)\right)$ bits
15:end for
16:Protocol at the server:
17:for $t+\tau\in\mathcal{T}_{T}$ do
18:  for $k\in\mathcal{K}$ do
19:   Run subroutine $\mathrm{DECODE}(\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}},s_{k},g,N)$ $\triangleright$ See Algorithm 3
20:  end for
21:  Compute $\hat{\mathbf{W}}_{t+\tau}\leftarrow\mathbf{W}_{t}+\sum_{k\in\mathcal{K}}p_{k}\hat{\mathbf{X}}_{t+\tau}^{k}$
22:  Set $\mathbf{W}_{t+\tau}\leftarrow\hat{\mathbf{W}}_{t+\tau}$ and broadcast $\mathbf{W}_{t+\tau}$ to all clients, or output $\mathbf{W}_{T}$ if $t+\tau=T$
23:end for
@@
1) Initialization: At the outset, client $k$ and the server agree on the privacy budget $\epsilon$ and privacy relaxation $\delta$ for $(\epsilon,\delta)$ DP (or $\epsilon$ for $\epsilon$-DP) in accordance with the system requirements, along with the parameters for LRSUQ as specified in Section III-B. The latter involves sharing a small random seed $s_{k}$ between client $k$ and the server to serve as a source of common randomness while fixing the lattice dimension $n$ and a lattice generator matrix $\mathbf{G}$. Both client $k$ and the server use the same random seed to initialize their respective PRNGs $\mathfrak{P}$ and $\tilde{\mathfrak{P}}$, ensuring that the outputs of the two PRNGs remain identical.
@@
Algorithm 2 $\mathrm{ENCODE}({\color[rgb]{0,0,0}\tilde{\mathbf{X}}_{t+\tau}^{k}},s_{k},(f,g),N)$
1:Inputs: Clipped model update vector $\tilde{\mathbf{X}}_{t+\tau}^{k}$, random seed $s_{k}$, pair of pdfs $(f,g)$, number of sub-vectors NN
@@
3:Partition $\tilde{\mathbf{X}}_{t+\tau}^{k}$ to $\{\tilde{\mathbf{X}}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$
4:for $j=1$ to NN do
5:  Sample $U_{t+\tau,j}^{k}\sim g$ by $\mathfrak{P}^{k}$ $\triangleright$ Initiated by seed $s_{k}$
6:  for $i=1,2,\ldots$ do $\triangleright$ Perform RSUQ
7:    Sample $\mathbf{V}_{t+\tau,j,i}^{k}\sim\mathrm{Unif}(\mathcal{P})$ by $\mathfrak{P}^{k}$ $\triangleright$ Initiated by seed $s_{k}$
8:   Find unique $\mathbf{M}_{t+\tau,j}^{k}\leftarrow Q_{\mathcal{P}}(\tilde{\mathbf{X}}_{t+\tau,j}^{k}/\beta(U_{t+\tau,j}^{k})-\mathbf{V}_{t+\tau,j,i}^{k})\in\mathbf{G}\mathbb{Z}^{n}$ $\triangleright$ Perform encoding
9:    Check if $\beta(U_{t+\tau,j}^{k})\cdot(\mathbf{M}_{t+\tau,j}^{k}+\mathbf{V}_{t+\tau,j,i}^{k})-\tilde{\mathbf{X}}_{t+\tau,j}^{k}\in L_{U_{t+\tau,j}^{k}}^{+}(f)$
10:   if Yes then $\triangleright$ Perform rejection sampling
11:     $H_{t+\tau,j}^{k}\leftarrow i$; Return $(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})$
12:   else
13:     Reject $i$ and repeat Step 7-9 with $i+1$
14:   end if
15:  end for
@@
17:return $\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}}$
Quantization: The $k$-th client clips the update $\mathbf{X}_{t+\tau}^{k}$ by a real number $\gamma>0$ to $\tilde{\mathbf{X}}_{t+\tau}^{k}$ (so that the $2$-norm of $\tilde{\mathbf{X}}_{t+\tau}^{k}$ is bounded by $\gamma$), and then divides $\tilde{\mathbf{X}}_{t+\tau}^{k}$ into $N:=\left\lceil\frac{m}{n}\right\rceil$ distinct $n$-dimensional sub-vectors $\tilde{\mathbf{X}}_{t+\tau,j}^{k}\in\mathbb{R}^{n}$, $j\in\{1,\ldots,N\}=:\mathcal{N}$.